OpenStudy (anonymous):

find all the critical points 3r/7r^2+8

OpenStudy (anonymous):

did you take the derivative?

OpenStudy (anonymous):

you should get \[\frac{-3(7x^2-8)}{(7x^2+8)^2}\]

OpenStudy (anonymous):

then set \[7x^2-8=0\] and solve in two steps \[7x^2=8\] \[x-\pm\sqrt{\frac{8}{7}}\]